OpenStudy (anonymous):

solve cot^2(x)=tan^2(x)

OpenStudy (aaronq):

\((cot(x))^2=(\dfrac{cosx}{sinx})^2\) \((cot(x))^2=(\dfrac{cosx}{sinx})^2=(tan(x))^2=(\dfrac{sinx}{cosx})^2\) (\(\dfrac{cosx}{sinx})^2=(\dfrac{sinx}{cosx})^2 \) only way this works is if sinx=cosx this occurs a \(\dfrac{\pi}4{}, \dfrac{3\pi}{4}, \dfrac{5\pi}{4}, etc\)